### v2

# TigerLLM - A Family of Bangla Large Language Models

Abstract
The development of Large Language Models (LLMs) remains heavily skewed towards English and a few other high-resource languages. This linguistic disparity is particularly evident for Bangla - the most spoken language. A few initiatives attempted to create open-source Bangla LLMs with performance still behind high-resource languages and limited reproducibility. To address this gap, we introduce TigerLLM - a family of Bangla LLMs. Our results demonstrate that these models surpass all open-source alternatives and also outperform larger proprietary models like GPT3.5 across standard benchmarks, establishing TigerLLM as the new baseline for future Bangla language modeling.
TigerLLM - A Family of Bangla Large Language Models
Nishat Raihan, Marcos Zampieri George Mason University, VA, USA mraihan2@gmu.edu
1 Introduction
LLMs have fundamentally transformed NLP by achieving exceptional performance across a broad range of tasks Brown et al. (2020); Raihan et al. (2025); Zhang et al. (2023). While these models exhibit unprecedented capabilities in language understanding, generation, reasoning, and specialized domains Raihan et al. (2024b), their advancements predominantly benefit high-resource languages Alam et al. (2024). This inequality is particularly noticeable for Bangla. Despite having about 237 million native speakers,111ethnologue.com/language/ben/ Bangla remains quite underserved in modern NLP advancements.
This under-representation stems primarily from the limitation of high-quality training data. While proprietary models like GPT-4 Brown et al. (2023) and Claude-3.5 Bai et al. (2024) demonstrate reasonable Bangla capabilities, open-source alternatives consistently underperform. Recent multilingual models such as Gemma-2 Team et al. (2024) and LLaMA 3.1 Dubey et al. (2024), despite leveraging diverse training corpora and advanced tokenization systems like TikTokenizer Corso et al. (2024), also fail to deliver satisfactory performance for Bangla.
| Base-LLM | Size | pt | corpora | ft | ft-dataset | Paper/Report? | Reproducibility? | |
| titu-Gemma | Gemma-2 | 2B | 4.4B | ✗ | ✗ | ✗ | ✗ | ✗ |
| titu-LLaMA | LLaMA-3.1 | 3B | 37B | ✗ | ✗ | ✗ | ✗ | ✗ |
| Bangla-LLaMA | LLaMA-3.2 | 3B | ✓ | ✗ | 172K | Orca-translated | ✓ | ✗ |
| G2B | Gemma-2 | 9B | ✗ | ✗ | 145K | Alpaca-translated | ✗ | ✗ |
| Bangla-LLaMA | LLaMA-2 | 13B | ✓ | ✗ | 145K | Alpaca-translated | ✗ | ✗ |
| TigerLLM | LLaMA-3.2 | 1B | 10M | Bangla-TextBook | 100K | Bangla-Instruct | ✓ | ✓ |
| TigerLLM | Gemma-2 | 9B | 10M | Bangla-TextBook | 100K | Bangla-Instruct | ✓ | ✓ |
1.1 Limitations of Bangla LLM Initiatives
Training
Recent attempts at developing Bangla LLMs (see Table 1) through continual pretraining titu-Gemma222huggingface.co/hishab/titulm-gemma-2-2b-v1.1 and model distillation approaches Zehady et al. (2024) have yielded low and non-reproducible results (see Table 2), often performing worse than their base models. The absence of technical documentation and academic publications further compounds this issue by making result reproduction impossible. Our investigation into these models’ performances reveals the need for improvement in the training process. While the unavailability of pretraining corpora limits our analysis of that phase, the finetuning approach demonstrates consistent problematic patterns.
Data
Most Bangla LLM initiatives rely on translated versions of synthetic datasets like Alpaca-Instruct Taori et al. (2023) and OpenOrca Mitra et al. (2023), which are generated through model distillation Hinton et al. (2015). This approach suffers from two fundamental limitations: (1) the datasets are generated by early GPT-3.5 Brown et al. (2020) releases, a model with limited Bangla support, resulting in suboptimal instruction quality, and (2) these English datasets are translated to Bangla using machine translation systems like Google Translate 333translate.google.com with limited quality checks, further degrading the training data quality. These cascading compromises in training data ultimately result in poor model performance.
1.2 Contributions
To address the recurring challenges in Bangla LLM development, we introduce three fundamental contributions:
1.
The Bangla-TextBook corpus, comprising 10 million tokens of carefully curated educational content across multiple domains, prioritizing content quality over scale.
2.
A high-quality Bangla-Instruct dataset of 100 thousand instruction-response pairs, generated through self-instruct Wang et al. (2023) and model distillation using state-of-the-art teacher models (GPT-4o and Claude-3.5-Sonnet).
3.
The Tiger-LLM family (1B and 9B parameters), featuring models pretrained and finetuned on our high-quality datasets, achieving 30-55% performance improvements over existing benchmarks.
All components are open-sourced to establish robust foundations for future Bangla language modeling research.444github.com/mraihan-gmu/TigerLLM/tree/main
2 Bangla-TextBook Corpus
Previous Bangla LLMs rely predominantly on corpora sourced from OSCAR Ortiz Suárez et al. (2020) and Common Crawl Bhattacharjee et al. (2022); Zehady et al. (2024), despite quality control challenges. While alternative Bangla corpora have emerged Bhattacharyya et al. (2023); Sketch Engine (2021), the absence of curated educational content remains a critical gap. This emphasis on data quality is particularly significant given recent findings by Gunasekar et al. (2023), which demonstrate that LLMs achieve superior performance through high-quality training data, even with reduced volume.
To bridge this gap, we present the Bangla-TextBook corpus, constructed exclusively from high-quality open-source educational materials published by the National Curriculum and Textbook Board of Bangladesh. We collect texts from 163 textbooks for Grades 6-12 resulting in a total of a total of 9,897,623 tokens and 697,903 sentences.
3 Bangla-Instruct
To address the limitations described in Section 1.1, we introduce Bangla-Instruct, a collection of 100,000 native Bangla instruction-response pairs bootstrapped using self-instruct Wang et al. (2023). While instruction datasets like Alpaca Taori et al. (2023) and OpenOrca Mitra et al. (2023) utilized GPT3 and GPT3.5 respectively, we significantly improve upon their approach by employing GPT-4 and Claude-3.5-Sonnet as our teacher models, leveraging their superior instruction-following capabilities.
Our dataset creation begins with 500 diverse seed tasks carefully curated by a team of 50 undergraduate and graduate students from leading Bangladeshi universities (Appendix A.1). These volunteers, spanning various academic disciplines and geographical regions of Bangladesh, ensure our seed tasks capture authentic linguistic patterns and cultural contexts. Each seed task undergoes multiple rounds of peer review to maintain quality and cultural sensitivity. Further information on quality control is presented in Appendix (Appendix A.3).
Our generation pipeline consists of four primary steps, each designed to maintain data quality and cultural authenticity (see Figure 1).
| MMLU-bn | PangBench-bn | BanglaQuaD | mHumanEval-bn | BEnQA | BanglaRQA | |
| understanding | multitasking | question answering | coding | knowledge | reasoning | |
| GPT3.5 | 0.55 | 0.55 | 0.50 | 0.56 | 0.50 | 0.49 |
| Gemini-Flash1.5 | 0.66 | 0.57 | 0.62 | 0.58 | 0.56 | 0.61 |
| GPT4o-mini | 0.67 | 0.62 | 0.65 | 0.56 | 0.60 | 0.60 |
| LLaMA3.2 (11B) | 0.22 | 0.19 | 0.21 | 0.15 | 0.18 | 0.20 |
| Gemma 2 (27B) | 0.35 | 0.51 | 0.43 | 0.64 | 0.50 | 0.56 |
| Pangea (7B) | 0.18 | 0.15 | 0.17 | 0.10 | 0.14 | 0.16 |
| Titu-LLM | 0.06 | 0.19 | 0.08 | 0.02 | 0.17 | 0.21 |
| Bong-LLaMA | 0.05 | 0.12 | 0.08 | 0.02 | 0.15 | 0.13 |
| Bangla-LLaMA | 0.02 | 0.08 | 0.05 | 0.10 | 0.11 | 0.09 |
| Bangla-Gemma | 0.18 | 0.15 | 0.12 | 0.10 | 0.22 | 0.19 |
| TigerLLM (1B) | 0.61 | 0.55 | 0.68 | 0.61 | 0.59 | 0.62 |
| TigerLLM (9B) | 0.72 | 0.68 | 0.70 | 0.63 | 0.65 | 0.68 |
1.
New Instruction Generation: Let be our seed task pool (Appendix A.2. For each generation round , we sample tasks uniformly at random from as demonstrations. Claude generates new instructions conditioned on these demonstrations while preserving Bangla linguistic patterns.
2.
Task Identification: Each new instruction undergoes classification by GPT-4o to determine task type . This classification helps in generating appropriate response formats.
3.
Instruction-Response Generation: For each instruction , we generate comprehensive responses using Claude that generates a candidate response.
4.
Filtering: Finally, GPT-4o carries out the multi-stage filtering process ensuring the Bengali language adherence, Cultural sensitivity, Response completeness and Discarding Similar instances (see Appendix A.3).
Valid instruction-response pairs are added back to the task pool, enabling iterative refinement. This process continues until we reach our target dataset size of 100K pairs, with each pair satisfying our quality criteria. The dual-model approach, combined with rigorous filtering, significantly reduces the likelihood of generating inappropriate or low-quality content, a common challenge in previous Bengali instruction datasets. More details are presented in Appendix A.
4 TigerLLM
We consider multiple candidate base models, similar to Raihan et al. (2024c)’s approach. 3 families of multilingual LLMs are chosen - LLaMA 3.2 (1B, 3B) Dubey et al. (2024), Gemma-2 (2B, 9B) Team et al. (2024) and Pangea (7B) Yue et al. (2024). Figure 2 depicts the final selection of model and a high-level overview of the process.
Continual Pretraining
We use the Bangla-TextBook corpus for the models to learn culture and language-specific nuances and gather sufficient and reliable knowledge from a set of high-quality texts. The pretraining phase has been carried out multiple times with empirical choices of hyper-parameters. the final set of hyperparameters is presented in Appendix B.
Model Distillation
We finetune our pretrained models using Bangla-Instruct. LoRA Hu et al. is not used, we implement full finetuning for better learning. To speed up the training process, we utilize Flash Attention Dao et al. (2022), we set key parameters: 2048 token maximum sequence length, batch size of 8, 4 gradient accumulation steps, and 3 epochs. Learning rate (), weight decay (0.02), and 10% warm-up steps ensure stable convergence. Table 5 in Appendix B lists complete hyperparameters.
5 Evaluation
Bangla LLM Benchmarks
Although there has been limited research on Bangla LLMs, several benchmarks have been established to assess their performance. We focus on five benchmarks specifically curated to evaluate Bangla LLMs across a diverse set tasks. For multitask understanding, we use the Bangla subset of MMLU-Pro Wang et al. (2024) and PangBench Yue et al. (2024). For question answering, we consider BanglaQuaD Rony et al. (2024), while for general knowledge, we use BEnQA Shafayat et al. (2024). For reasoning tasks, we refer to BanglaRQA Ekram et al. (2022), and for coding tasks, we utilize the Bangla subset of mHumanEval Raihan et al. (2024a).
Results
We present the results obtained by the two TigerLLM models compared to a variety of strong LLM baselines in Table 2. The performance comparison of various models on Bangla-specific benchmarks reveals a common trend. The fine-tuned models generally perform worse than their base counterparts across most tasks. In particular, the results reported by the authors are not reproducible, as mentioned in Section 1.1. However, TigerLLM is the only finetuned model, consistently outperforming both its base and fine-tuned variants across all tasks. Even the 1B variant does better than most models, falling short to only its 9B counterpart, further validating our emphasis on high-quality data (Section 3).
Takeaways
TigerLLM demonstrates that carefully curated, high-quality datasets can yield superior performance even with smaller model sizes. Our results show that the 1B parameter model outperforms larger alternatives across multiple benchmarks, emphasizing the importance of data quality over quantity. The success of our Bangla-TextBook corpus and Bangla-Instruct dataset establishes a new paradigm for low-resource language model development.
6 Conclusion and Future Work
This paper introduces TigerLLM, a family of state-of-the-art Bangla language models that outperforms existing alternatives across six benchmarks. TigerLLM’s success stems from two key innovations: (1) the high-quality Bangla-TextBook corpus derived from educational materials, and (2) the carefully curated Bangla-Instruct dataset generated using advanced teacher models. The three resources introduced here (corpus, instruct dataset, and models) establish a robust foundation for future Bangla language modeling research. Together they will contribute to speed up advances in Bangla language modeling.
Future work will focus on (1) qualitative analysis of model performance, (2) expanding the corpus to include more diverse domains, (3) scaling up model sizes while maintaining quality, and (4) developing more sophisticated evaluation metrics specific to Bangla language tasks.
Limitations
While TigerLLM demonstrates impressive performance, several limitations warrant acknowledgment. First, our Bangla-TextBook corpus, though carefully curated, is limited to educational materials from grades 6-12, potentially missing broader linguistic patterns present in other domains. The 10 million token size, while sufficient for our current models, may constrain scaling to larger architectures. Additionally, our Bangla-Instruct dataset, despite its quality-focused generation process, covers only a subset of possible instruction types and may not fully capture the complexity of real-world Bangla language use cases.
Furthermore, our models are currently limited to 1B and 9B parameters, primarily due to computational constraints and our emphasis on thorough experimentation with smaller computationally efficient architectures. While this approach enabled rapid iteration and quality-focused development, it may not fully exploit the potential benefits of larger model scales.
Ethical Considerations
Our work prioritizes ethical considerations throughout the development process. The Bangla-TextBook corpus uses open-source publicly available educational materials from the National Curriculum and Textbook Board of Bangladesh. The volunteer-driven seed task creation process incorporated diverse perspectives while maintaining cultural sensitivity and avoiding harmful biases.
We implemented rigorous filtering mechanisms to ensure cultural appropriateness, gender neutrality, and religious sensitivity in our instruction dataset. The multi-stage review process, involving both automated checks and human verification, helps prevent the propagation of harmful stereotypes or biases. Additionally, our open-source approach promotes transparency and enables community oversight of model behavior.
We strongly recommend that users implement appropriate safeguards when deploying TigerLLM in production environments, particularly for applications involving sensitive information or critical decision-making.
Appendix A Bangla-Instruct Curation
A.1 Volunteer Information
The seed tasks were created by 50 undergraduate and graduate students from various universities across Bangladesh, ensuring geographical and academic diversity:
15 students from Computer Science and Engineering.
10 students from Bengali Literature.
10 students from Business Administration.
8 students from Science and Engineering.
7 students from Social Sciences.
Each volunteer contributed 10 diverse instructions, resulting in our initial pool of 500 seed tasks. The distribution ensured coverage across multiple domains while preserving authentic Bengali linguistic patterns and cultural contexts.
A.2 The Seed Dataset
Our seed dataset comprises 10 distinct categories, carefully chosen to cover a broad spectrum of tasks relevant to Bengali language and culture:
1.
Cultural Knowledge and Heritage (): Tasks focusing on Bengali traditions, festivals, folk tales, and historical events. These include explaining cultural practices, describing traditional ceremonies, and discussing historical significance of various customs.
2.
Academic Writing (): Structured writing tasks ranging from essay outlines to full academic compositions. Topics cover various academic disciplines while maintaining Bengali writing conventions and scholarly standards.
3.
Mathematical Problem Solving (): Tasks involving mathematical concepts explained in Bengali, including algebra, geometry, and arithmetic. Special attention is given to Bengali mathematical terminology and local problem-solving contexts.
4.
Programming and Technical (): Programming problems described in Bengali with solutions in standard programming languages. Includes algorithm explanation, code documentation, and technical concept elaboration in Bengali.
5.
Creative Writing (): Open-ended creative tasks including story writing, poetry composition, and descriptive passages. Emphasizes Bengali literary devices, metaphors, and cultural storytelling elements.
6.
Scientific Explanation (): Tasks requiring clear explanation of scientific concepts in Bengali, focusing on making complex ideas accessible while maintaining technical accuracy. Covers physics, chemistry, biology, and environmental science.
7.
Business and Economics (): Professional writing tasks including business case analyses, market reports, and economic concept explanations. Incorporates local business contexts and Bengali business terminology.
8.
Social Issues Analysis (): Critical analysis tasks addressing contemporary social issues in Bangladesh and Bengali society. Includes problem identification, cause analysis, and solution proposition.
9.
Data Analysis and Statistics (): Tasks involving interpretation and analysis of data presented in Bengali, including statistical concepts explanation, data visualization description, and numerical analysis.
10.
Language and Translation (): Tasks focused on Bengali language mastery, including idiom explanation, translation between Bengali and English, and linguistic analysis of Bengali texts.
Each category accounts for approximately 10% of the seed dataset ( tasks per category), ensuring balanced representation across domains. The tasks within each category vary in complexity level: 40% basic, 40% intermediate, and 20% advanced, based on linguistic complexity and cognitive demand.
A.3 Filtering Methodology
Our filtering process implements the following criteria:
1.
Language Adherence ()
Bengali Word Ratio:
Unicode Consistency:
Grammar Check: Using GPT-4o’s Bengali grammar scoring function
2.
Cultural Sensitivity ()
Religious Neutrality: on our bias scale
Regional Inclusivity: No specific region/dialect preference
Gender Representation: Balanced pronouns and roles
Political Neutrality: Avoidance of partisan content
3.
Content Quality ()
Minimum Length: where is task type
Coherence Score: between instruction and response
Factual Accuracy: Verified against Bengali Wikipedia
Format Adherence: Proper paragraph breaks, lists, or code blocks
4.
Novelty Verification ()
Similarity Threshold:
Lexical Diversity: Minimum Type-Token Ratio of 0.4
Response Uniqueness: No duplicate responses within same category
Task Format Variation: Ensure uniform distribution across formats
A pair is accepted if and only if:
This rigorous filtering ensures the quality and diversity of our final dataset while maintaining Bengali linguistic and cultural authenticity.
Appendix B Experimentation Details
B.1 Experimental Setup
Pretraining
We utilize a Lambda Labs555lambdalabs.com cluster with 8 NVIDIA A100 GPUs (40GB each), 512GB RAM, and 2TB storage. The distributed training setup enables efficient parallel processing, completing the pretraining in approximately 120 hours on this high-performance configuration with gradient checkpointing enabled.
Finetuning
We conduct finetuning on a single NVIDIA A100 (40GB) through Google Colab666colab.research.google.com, supported by 80GB RAM and 256GB storage. The process completes in approximately 96 hours, proving sufficient for model adaptation and task-specific optimization with minimal computational overhead.
B.2 Pretraining HyperParameters
| Hyperparameter | Value |
| Per device train batch size | 64 |
| Gradient accumulation steps | 16 |
| Number of training epochs | 4 |
| Learning rate | |
| FP16 | False |
| BF16 | True |
| Dataloader num workers | 8 |
| Gradient checkpointing | True |
| Logging steps | 1000 |
| DDP find unused parameters | False |
| Max gradient norm | 1.0 |
| Warmup steps | 1000 |
| Evaluation strategy | steps |
| Evaluation steps | 1,000 |
| Save strategy | steps |
| Save steps | 1,000 |
| Save total limit | 3 |
| Load best model at end | True |
| Metric for best model | loss |
| Greater is better | False |
B.3 Finetuning Hyperparameters
| Hyperparameter | Value |
| Max Sequence Length | 2048 |
| Batch Size (Train/Eval) | 16 |
| Gradient Accumulation Steps | 4 |
| Number of Epochs | 3 |
| Learning Rate | 1e-5 |
| Weight Decay | 0.02 |
| Warmup Steps | 10% |
| Optimizer | AdamW (8-bit) |
| LR Scheduler | Cosine |
| Precision | BF16 |
| Evaluation Strategy | Steps |
| Evaluation Steps | 50 |
| Save Strategy | Steps |
| Save Steps | Varies |
| Seed | 42 |
| Hyperparameter | Value |
| Max Sequence Length | 2048 |
| Batch Size (Train/Eval) | 32 |
| Gradient Accumulation Steps | 8 |
| Number of Epochs | 3 |
| Learning Rate | 1e-6 |
| Weight Decay | 0.04 |
| Warmup Steps | 15% |
| Optimizer | AdamW (8-bit) |
| LR Scheduler | Cosine |
| Precision | BF16 |
| Evaluation Strategy | Steps |
| Evaluation Steps | 250 |
| Save Strategy | Steps |
| Save Steps | Varies |
| Seed | 42 |
Appendix C TigerLLM - Training Pipeline
Figure 2 illustrates the multi-stage training pipeline for producing both TigerLLM (1B) and TigerLLM (9B). It begins with two pre-trained language models—LLaMA 3.2 (1B) and Gemma 2 (9B)—chosen for their robust foundational capacities. These models then undergo continual pretraining (see Figure 3) on a specialized Bangla-TextBook corpus, which infuses them with a richer understanding of the Bangla language, including its context-specific nuances, stylistic variations, and domain-specific terminology.
Following this continual pretraining step (Figure 3), the models are finetuned on a carefully curated Bangla-Instruct dataset. This additional phase (Figure 4) ensures that the trained models can effectively process and respond to Bangla instructions, yielding outputs that are more coherent, context-aware, and aligned with users’ informational needs.
In essence, by blending the foundational strengths of LLaMA and Gemma with specialized Bangla corpora and instruction-oriented finetuning, the final TigerLLM models emerge as optimized solutions capable of delivering high-quality, instruction-following responses tailored to Bangla-language tasks.

### v3

@@
 TigerLLM - A Family of Bangla Large Language Models
-Nishat Raihan, Marcos Zampieri George Mason University, VA, USA mraihan2@gmu.edu
+Nishat Raihan George Mason University Fairfax, VA, USA mraihan2@gmu.edu Marcos Zampieri George Mason University Fairfax, VA, USA mzampier@gmu.edu
 1 Introduction
 LLMs have fundamentally transformed NLP by achieving exceptional performance across a broad range of tasks Brown et al.
+(2020); Chowdhery et al.
 (2022); Raihan et al.
-(2025); Zhang et al.
-(2023).
-While these models exhibit unprecedented capabilities in language understanding, generation, reasoning, and specialized domains Raihan et al.
-(2024b), their advancements predominantly benefit high-resource languages Alam et al.
+(2025c).
+While these models exhibit unprecedented capabilities in language understanding, generation, reasoning, and specialized applications, their advancements predominantly benefit high-resource languages Alam et al.
 (2024).
@@
 (2024) demonstrate reasonable Bangla capabilities, open-source alternatives consistently underperform.
-Recent multilingual models such as Gemma-2 Team et al.
+Recent multilingual models such as Gemma-2 Gemma et al.
 (2024) and LLaMA 3.1 Dubey et al.
@@
 Training
-Recent attempts at developing Bangla LLMs (see Table 1) through continual pretraining titu-Gemma222huggingface.co/hishab/titulm-gemma-2-2b-v1.1 and model distillation approaches Zehady et al.
+Recent attempts at developing Bangla LLMs (see Table 1) through continual pretraining (titu-Gemma) and model distillation approaches Zehady et al.
 (2024) have yielded low and non-reproducible results (see Table 2), often performing worse than their base models.
@@
 This approach suffers from two fundamental limitations: (1) the datasets are generated by early GPT-3.5 Brown et al.
-(2020) releases, a model with limited Bangla support, resulting in suboptimal instruction quality, and (2) these English datasets are translated to Bangla using machine translation systems like Google Translate 333translate.google.com with limited quality checks, further degrading the training data quality.
+(2020) releases, a model with limited Bangla support, resulting in suboptimal instruction quality, and (2) these English datasets are translated to Bangla using machine translation systems like Google Translate with limited quality checks, further degrading the training data quality.
 These cascading compromises in training data ultimately result in poor model performance.
@@
 The Tiger-LLM family (1B and 9B parameters), featuring models pretrained and finetuned on our high-quality datasets, achieving 30-55% performance improvements over existing benchmarks.
-All components are open-sourced to establish robust foundations for future Bangla language modeling research.444github.com/mraihan-gmu/TigerLLM/tree/main
+All components are open-sourced to establish robust foundations for future Bangla language modeling research.222https://github.com/mraihan-gmu/TigerLLM/tree/main/
+2 Related Work
+Early transformer-based encoder-only pre-trained language models such as BERT Devlin et al.
+(2019) concentrate on high-resource languages like English.
+Subsequent work adapts them to mid- and low-resource contexts through continued pre-training and task-specific finetuning.
+In Bangla, for instance, Sami et al.
+(2022) present BanglaBERT, demonstrating that a dedicated monolingual encoder markedly improves downstream classification and QA relative to multilingual baselines.
+The shift to decoder-only models has produced large multilingual models — e.g.
+BLOOM Le Scao et al.
+(2022), Llama3 Dubey et al.
+(2024), and Aya Üstün et al.
+(2024)—that cover dozens of under-represented languages.
+Yet empirical analyses reveal that these models still perform best when prompted in high-resource languages, with significant degradation for languages such as Bangla or Swahili Raihan et al.
+(2025a); Jin et al.
+(2024).
+As discussed in the previous section, dedicated Bangla decoder models remain scarce and fragmented.
+GPT2-Bangla Bhattacharjee et al.
+(2023) continues GPT-2 pre-training on a 4GB Bangla corpus, while Bong-LLaMA Zehady et al.
+(2024) and the titu-Gemma333https://huggingface.co/hishab/titulm-gemma-2-2b-v1.1 checkpoint attempt instruction tuning on translated datasets.
+These efforts often lack rigorous evaluation protocols, transparent data curation, or reproducible training pipelines—as reflected in the inconsistent results summarized in Table 1.
+Consequently, a clear methodological gap persists in developing open, reproducible decoder-only LLMs that natively support Bangla and other low-resource languages.
 3 Bangla-TextBook Corpus
 Previous Bangla LLMs rely predominantly on corpora sourced from OSCAR Ortiz Suárez et al.
-(2020) and Common Crawl Bhattacharjee et al.
+and Common Crawl Bhattacharjee et al.
 (2022); Zehady et al.
@@
 While alternative Bangla corpora have emerged Bhattacharyya et al.
-(2023); Sketch Engine (2021), the absence of curated educational content remains a critical gap.
+(2023), the absence of curated educational content remains a critical gap.
 This emphasis on data quality is particularly significant given recent findings by Gunasekar et al.
-(2023), which demonstrate that LLMs achieve superior performance through high-quality training data, even with reduced volume.
+(2023) and Raihan et al.
+(2025b), which demonstrate that LLMs achieve superior performance through high-quality training data, even with reduced volume.
 To bridge this gap, we present the Bangla-TextBook corpus, constructed exclusively from high-quality open-source educational materials published by the National Curriculum and Textbook Board of Bangladesh.
-We collect texts from 163 textbooks for Grades 6-12 resulting in a total of a total of 9,897,623 tokens and 697,903 sentences.
+We collect texts from 163 textbooks for Grades 6-12, resulting in a total of 9,897,623 tokens and 697,903 sentences.
 4 Bangla-Instruct
@@
 Our generation pipeline consists of four primary steps, each designed to maintain data quality and cultural authenticity (see Figure 1).
+(1) Seed & Instruction Generation: We begin with a human-curated seed pool drawn from 50 volunteers representing five academic disciplines across Bangladesh (see Appendix A.1).
+At every generation round , we sample seed tasks and prompt Claude to create a candidate batch of instructions , expanding coverage of the ten seed categories listed in Appendix A.2 while preserving authentic linguistic patterns.
+(2) Task Typing: Each instruction is classified by GPT-4o into , providing the expected answer style and the minimum-length threshold used in subsequent filtering.
+(3) Response Drafting: Conditioned on , Claude produces a comprehensive response . We retain the highest-scoring draft according to an internal coherence metric .
+(4) Multi-stage Filtering: GPT-4o applies the four-criteria filter —Language (), Cultural (), Quality (), and Novelty () (see Appendix A.3).
+On average, ~63% of pairs pass , yielding a balanced complexity mix (40% basic, 40% intermediate, 20% advanced).
+Valid pairs are appended to , and the loop continues until 100K high-quality instruction–response pairs are reached.
+By coupling two complementary LLMs with strict verification and a human-seeded, domain-balanced task pool, our pipeline mitigates error propagation and preserves cultural nuance—addressing shortcomings observed in earlier Bengali instruction datasets (see Appendix A for full statistics).
+5 TigerLLM
+As candidate base models, we consider 3 families of multilingual LLMs - LLaMA 3.2 (1B, 3B) Dubey et al. (2024), Gemma-2 (2B, 9B) Gemma et al. (2024) and Pangea (7B) Yue et al. (2024).
+Evolution of TigerLLM
+Figure 2 depicts the final selection of the models and a high-level overview of the process.
+Upon the selection phase, we finalize two pre-trained language models—LLaMA 3.2 (1B) and Gemma 2 (9B)—chosen for their robust foundational capacities.
+These models then undergo continual pretraining (see Figure 3) on a specialized Bangla-TextBook corpus, which infuses them with a richer understanding of the Bangla language, including its context-specific nuances, stylistic variations, and domain-specific terminology.
+Pretraining
+We utilize a computing cluster with 8 NVIDIA A100 GPUs (40GB each), 512GB RAM, and 2TB storage.
+The distributed training setup enables efficient parallel processing, completing the pretraining in approximately 120 hours on this high-performance configuration with gradient checkpointing enabled.
+Continual Pretraining
+We use the Bangla-TextBook corpus for the models to learn culture and language-specific nuances and gather sufficient and reliable knowledge from a set of high-quality texts.
+The pretraining phase has been carried out multiple times with empirical choices of hyper-parameters.
+Finetuning
+We conduct finetuning on a single NVIDIA A100 (40GB) through Google Colab444colab.research.google.com, supported by 80GB RAM and 256GB storage.
+The process completes in approximately 96 hours, proving sufficient for model adaptation and task-specific optimization with minimal computational overhead.
+Model Distillation
+Following this continual pretraining step, the models are finetuned on a carefully curated Bangla-Instruct dataset (Figure 4).
+LoRA Hu et al.
+(2021) is not used, we implement full finetuning for better learning.
+To speed up the training process, we utilize Flash Attention Dao et al.
+(2022), we set key parameters: 2048 token maximum sequence length, batch size of 8, 4 gradient accumulation steps, and 3 epochs.
+Learning rate (), weight decay (0.02), and 10% warm-up steps ensure stable convergence.
+Table 5 in Appendix B lists complete hyperparameters.
 | MMLU-bn | PangBench-bn | BanglaQuaD | mHumanEval-bn | BEnQA | BanglaRQA | |
@@
 | TigerLLM (9B) | 0.72 | 0.68 | 0.70 | 0.63 | 0.65 | 0.68 |
-1.
-New Instruction Generation: Let be our seed task pool (Appendix A.2.
-For each generation round , we sample tasks uniformly at random from as demonstrations.
-Claude generates new instructions conditioned on these demonstrations while preserving Bangla linguistic patterns.
-2.
-Task Identification: Each new instruction undergoes classification by GPT-4o to determine task type . This classification helps in generating appropriate response formats.
-3.
-Instruction-Response Generation: For each instruction , we generate comprehensive responses using Claude that generates a candidate response.
-4.
-Filtering: Finally, GPT-4o carries out the multi-stage filtering process ensuring the Bengali language adherence, Cultural sensitivity, Response completeness and Discarding Similar instances (see Appendix A.3).
-Valid instruction-response pairs are added back to the task pool, enabling iterative refinement.
-This process continues until we reach our target dataset size of 100K pairs, with each pair satisfying our quality criteria.
-The dual-model approach, combined with rigorous filtering, significantly reduces the likelihood of generating inappropriate or low-quality content, a common challenge in previous Bengali instruction datasets.
-More details are presented in Appendix A.
-4 TigerLLM
-We consider multiple candidate base models, similar to Raihan et al.
-(2024c)’s approach.
-3 families of multilingual LLMs are chosen - LLaMA 3.2 (1B, 3B) Dubey et al.
-(2024), Gemma-2 (2B, 9B) Team et al.
-(2024) and Pangea (7B) Yue et al.
-(2024).
-Figure 2 depicts the final selection of model and a high-level overview of the process.
-Continual Pretraining
-We use the Bangla-TextBook corpus for the models to learn culture and language-specific nuances and gather sufficient and reliable knowledge from a set of high-quality texts.
-The pretraining phase has been carried out multiple times with empirical choices of hyper-parameters.
-the final set of hyperparameters is presented in Appendix B.
-Model Distillation
-We finetune our pretrained models using Bangla-Instruct.
-LoRA Hu et al.
-is not used, we implement full finetuning for better learning.
-To speed up the training process, we utilize Flash Attention Dao et al.
-(2022), we set key parameters: 2048 token maximum sequence length, batch size of 8, 4 gradient accumulation steps, and 3 epochs.
-Learning rate (), weight decay (0.02), and 10% warm-up steps ensure stable convergence.
-Table 5 in Appendix B lists complete hyperparameters.
+By blending the foundational strengths of LLaMA and Gemma with specialized Bangla corpora and instruction-oriented finetuning, the final TigerLLM models emerge as optimized solutions capable of delivering high-quality, instruction-following responses tailored to Bangla-language tasks.
 6 Evaluation
@@
 For reasoning tasks, we refer to BanglaRQA Ekram et al.
-(2022), and for coding tasks, we utilize the Bangla subset of mHumanEval Raihan et al.
-(2024a).
+(2022).
 Results
@@
 This paper introduces TigerLLM, a family of state-of-the-art Bangla language models that outperforms existing alternatives across six benchmarks.
-TigerLLM’s success stems from two key innovations: (1) the high-quality Bangla-TextBook corpus derived from educational materials, and (2) the carefully curated Bangla-Instruct dataset generated using advanced teacher models.
-The three resources introduced here (corpus, instruct dataset, and models) establish a robust foundation for future Bangla language modeling research.
-Together they will contribute to speed up advances in Bangla language modeling.
-Future work will focus on (1) qualitative analysis of model performance, (2) expanding the corpus to include more diverse domains, (3) scaling up model sizes while maintaining quality, and (4) developing more sophisticated evaluation metrics specific to Bangla language tasks.
+TigerLLM’s success stems from two key innovations: (1) the high-quality Bangla-TextBook corpus derived from educational materials and (2) the carefully curated Bangla-Instruct dataset generated using advanced teacher models.
+The three resources introduced here (corpus, instruction dataset, and models) establish a robust foundation for future Bangla language modeling research.
+Together, they will contribute to speeding up advances in Bangla language modeling.
+In future work we will conduct a deeper qualitative analysis of the model’s behavior, broaden the corpus to cover a wider array of domains, scale the model to larger parameter counts without compromising quality, and devise richer evaluation metrics tailored specifically to Bangla tasks.
 Limitations
-While TigerLLM demonstrates impressive performance, several limitations warrant acknowledgment.
+While TigerLLM delivers state-of-the-art performance, several limitations warrant acknowledgment.
 First, our Bangla-TextBook corpus, though carefully curated, is limited to educational materials from grades 6-12, potentially missing broader linguistic patterns present in other domains.
@@
 Appendix B Experimentation Details
-B.1 Experimental Setup
-Pretraining
-We utilize a Lambda Labs555lambdalabs.com cluster with 8 NVIDIA A100 GPUs (40GB each), 512GB RAM, and 2TB storage.
-The distributed training setup enables efficient parallel processing, completing the pretraining in approximately 120 hours on this high-performance configuration with gradient checkpointing enabled.
-Finetuning
-We conduct finetuning on a single NVIDIA A100 (40GB) through Google Colab666colab.research.google.com, supported by 80GB RAM and 256GB storage.
-The process completes in approximately 96 hours, proving sufficient for model adaptation and task-specific optimization with minimal computational overhead.
-B.2 Pretraining HyperParameters
+B.1 Pretraining HyperParameters
 | Hyperparameter | Value |
@@
 | Greater is better | False |
-B.3 Finetuning Hyperparameters
-| Hyperparameter | Value |
+B.2 Finetuning Hyperparameters
+| Parameter | Value |
 | Max Sequence Length | 2048 |
@@
 | Seed | 42 |
-| Hyperparameter | Value |
+| Parameter | Value |
 | Max Sequence Length | 2048 |
@@
 | Seed | 42 |
-Appendix C TigerLLM - Training Pipeline
-Figure 2 illustrates the multi-stage training pipeline for producing both TigerLLM (1B) and TigerLLM (9B).
-It begins with two pre-trained language models—LLaMA 3.2 (1B) and Gemma 2 (9B)—chosen for their robust foundational capacities.
-These models then undergo continual pretraining (see Figure 3) on a specialized Bangla-TextBook corpus, which infuses them with a richer understanding of the Bangla language, including its context-specific nuances, stylistic variations, and domain-specific terminology.
-Following this continual pretraining step (Figure 3), the models are finetuned on a carefully curated Bangla-Instruct dataset.
-This additional phase (Figure 4) ensures that the trained models can effectively process and respond to Bangla instructions, yielding outputs that are more coherent, context-aware, and aligned with users’ informational needs.
-In essence, by blending the foundational strengths of LLaMA and Gemma with specialized Bangla corpora and instruction-oriented finetuning, the final TigerLLM models emerge as optimized solutions capable of delivering high-quality, instruction-following responses tailored to Bangla-language tasks.